4x-2/5y=-6 4x+3/5y=-1 solve the system by the addition method
Answer by checkley77(12844)

4x-2/5y=-6 4x+3/5y=-1 subtract.
-------------------
-2/5y-3/5y=-5
-5/5y=-5
-y=-5
y=5 ans.
4x-2/5*5=-6
4x-2=-6
4x=-6+2
4x=-4
x=-1 ans.
Proof:
4*-1+3/5*5=-1
-4+3=-1
-1=-1